The position of a SHO as a function of time is given by x = 7.1cos 7πt  + π 

                          3      6

where t is in seconds and x is in centimetres. Find:

a)               the period and frequency expressed as fractions                        

b)              the velocity at t = 2.5 (to 2 sig. fig.)                                            

 


1
Expert's answer
2021-11-29T11:45:16-0500

"x=7.1\cos(7\pi t+\pi/36)". So, we have


a) "T=2\pi/\omega=2\pi/(7\pi)=2/7\ (s)"


"f=1/T=1/(2/7)=7/2\ (Hz)"


b) "v=dx/dt=-1.56\cdot\sin(7\pi t+\pi/36)"


"v(2.5)=-1.56\cdot\sin(7\pi \cdot2.5+\pi/36)=-1.6\ (m/s)"